I have right triangle. Other angles are 38 and 52. The longest side is 4.25 inches. What length should the other two sides be?
Answer by stanbon(75887)

I have right triangle. Other angles are 38 and 52. The longest side is 4.25 inches. What length should the other two sides be?
-----
3rd angle = (180)-(38+52) = 90 degrees
------
So the hypotenuse = 4.25
-------
Draw the picture.
Side opposite the 38 degree angle = 4.25*sin(38) = 2.62
----
Side opposite the 52 degree angle = 4.25sin(52) = 3.35
============================
Cheers,
Stan H.
=========